OpenStudy (anonymous):

one circle has a radius of 8 m, and a second circle has a radius of 10 m. What is the difference in area of the two circles? Use pi = 3.14 as an approximate

OpenStudy (anonymous):

Area of a circle = pir^2 small circle = pi(8)^2=64pi large circle=pi(10)^2=100pi now subtract them and use pi=3.14

OpenStudy (anonymous):

\[ pir^2 \pi(8)^2=64\pi \\pi(10)^2=100\pi\pi=3.14\]

OpenStudy (anonymous):

hey can u rewrite that using the "Equation" box? XD

OpenStudy (anonymous):

ok :)small circle \[= \pi(8)^2 = 64\pi \] large \[\ = 10^2pi=100\pi\] now subract them

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

do i subbtract 100(pi) - 64(pi)?

OpenStudy (anonymous):

\[100(\pi) - 64(\pi)\]

OpenStudy (anonymous):

i think my answer is 36(pi)